# Reveal-Bangla: A Dataset for Cross-Lingual Multi-Step Reasoning Evaluation

### v2

@@ 1 Introduction
Large Language Models (LLMs) have demonstrated remarkable versatility across a wide spectrum of natural language processing tasks (Radford et al., 2019). A pivotal breakthrough in enhancing their complex reasoning capabilities has been the introduction of Chain-of-Thought (CoT) prompting (Wei et al., 2022), which encourages models to generate intermediate reasoning steps before arriving at final answers, yielding substantial performance improvements (White et al., 2024; Wang et al., 2022). Despite these advances, the evaluation of LLM reasoning capabilities remains heavily skewed toward high-resource languages, creating significant gaps in our understanding of how these models perform across linguistically diverse contexts. In this work, we focus specifically on the Bangla language, which boasts 268 million speakers and ranks as the sixth most spoken language globally [^2] https://en.wikipedia.org/wiki/List_of_languages_by_total_number_of_speakers. As the native language of Bangladesh and the second most prominent Indo-Aryan language after Hindi Eberhard et al. (2021), Bangla represents a critical case study for cross-lingual reasoning evaluation. The growing technological transformation in densely populated and economically emerging regions where Bangla is spoken Rahman (2024) underscores the urgent need for developing faithful AI technologies that can enhance social welfare and economic opportunities. While recent work in Bangla focused on simple extractive or multiple-choice question answering Ekram et al. (2022); Shafayat et al. (2024); Rony et al. (2024), to our knowledge, no datasets with human-validated reasoning steps are available for this language. This lack of resources hinders our ability to assess and improve the reasoning capabilities of LLMs in the Bangla language.
@@
In this work, we address this gap by introducing Reveal-Bangla, a manually translated Bangla version of a subset of the English Reveal dataset, containing annotated multi-step reasoning chains with gold answers. We exploit our resource and its original English counterpart to evaluate the abilities of two small language models—both proficient in Bangla and English, but one predominantly English-centric, and the other mainly Bangla-centric—in exploiting reasoning step to produce the correct answers given a query, following recent work showing how non-English languages can harm reasoning abilities in LLMs (Qi et al., 2025).
@@
Moreover, recent cross-lingual studies have revealed that generated reasoning chains often exhibit inconsistencies and produce misleading intermediate steps, raising questions about their explanatory reliability (Lanham et al., 2023; Paul et al., 2024). To address these concerns, post-hoc attribution techniques have emerged as valuable tools for analyzing models’ internal processes by assigning importance scores to context elements such as retrieved documents (Qi et al., 2024; Cohen-Wang et al., 2024) and reasoning chains (Cohen-Wang et al., 2025), thereby revealing their contribution to final predictions. We exploit a similar methodology using the ContextCite method (Cohen-Wang et al., 2024) to examine how reasoning steps contribute to model answers in English and Bangla, highlighting different patterns of importance across the two languages.
@@ 2 Related Work
Large Language Models (LLMs) operate as probabilistic sequence predictors, estimating the likelihood of the next token given previous context Vaswani et al. (2017); Radford et al. (2019). For practical application, explicit training on instructions was found to further improve answer quality (Sanh et al., 2022; Wang et al., 2022). Recently, eliciting reasoning from LLMs, e.g. via step-by-step Chain of Thought reasoning (CoT, Wei et al., 2022), was found to further improve the response accuracy for complex queries.
Some popular reasoning datasets in English include StrategyQA (Geva et al., 2021), featuring reasoning- and knowledge-intensive yes/no queries; Fermi (Kalyan et al., 2021), comprising estimation questions that require numerical answers and a blend of knowledge and reasoning; MuSiQue Trivedi et al. (2022), which includes multi-hop reasoning questions with free-text entity answers, generated from Wikipedia paragraphs; and Sports Understanding Srivastava et al. (2022), consisting of yes/no questions that demand reasoning about sports players, leagues, and maneuvers. Jacovi et al. (2024) combined the aforementioned datasets and human-annotated each LLM-generated step in terms of attribution relative to provided Wikipedia paragraphs and logical coherence in light of previous reasoning steps. The resulting dataset, dubbed Reveal, was used to prompt capable LLMs in a chain-of-thought setting and analyzed using Natural Language Inference (NLI) classifiers to evaluate model-generated responses. In our work, we manually translate a subset of Reveal into Bangla and adopt their setup to evaluate cross-lingual English-Bangla models.
Recently, Jin et al. (2024) explored small and large parameterized models, revealing a linear relationship between accuracy and the number of reasoning steps. We conduct a similar analysis, focusing particularly on small language models (SLMs) with a manageable size (1B parameters). SLMs were recently found capable of high-quality answers in RAG setups (Huang et al., 2024; Liu et al., 2024b), with relevant input information compensating for their limited reasoning abilities. In this work, we use annotated CoTs produced by larger LLMs to investigate whether SLMs can effectively leverage reasoning information in English and Bangla.
@@ 3.1 Data Collection
We start by selecting a subset of the Reveal dataset. [^3] https://huggingface.co/datasets/google/reveal Provided we want to test the ability of SLMs to obtain the correct answer given valid reasoning chains, we focus specifically on examples having all reasoning steps as either logical or fully attributable to the provided Wikipedia paragraphs. Furthermore, among the three models considered by the Reveal authors to generate answers, we decided to choose the two models with the most answers, i.e., Flan-UL2-20B Tay et al. (2022) and GPT-3 (text-davinci-003, Brown et al., 2020). [^4] We do not include Flan-PaLM-540B Longpre et al. (2023) due to our limited evaluation resources.
@@ 4 Evaluation
For our evaluation on Reveal and our Bangla variant, we use Llama-3.2-1B-Instruct (or EngLlama) Grattafiori et al. (2024) and BanglaLLama-3.2-1b-bangla-alpaca-orca-instruct-v0.0.1 (or BenLlama) Zehady et al. (2024). EngLlama is a popular English-centric multilingual SLM, while BenLlama is a Bangla-centric model fine-tuned from EngLlama using Bangla-Alpaca-Orca, a collection of instruction tuning examples including the popular Alpaca and OpenOrca datasets (Taori et al., 2023; Lian et al., 2023) automatically translated into Bangla. Importantly, despite their different language focus, both models maintain answering capabilities in both English and Bangla, motivating our cross-lingual analysis.
@@
To verify the accuracy of the model-generated final answer against the actual final answer, we choose mDeBERTa-v3-base-xnli-multilingual-nli-2mil7 Laurer et al. (2022) model as it is the only NLI model that supports both English and Bangla. We consider entailment labels as correct answers and contradict as otherwise. As this NLI model additionally verdicts neutral, authors manually verify the response to classify it as valid or not. Furthermore, as language detection tools such as langdetect Shuyo (2010) do not support Bangla, we manually assign contradict to answers generated in scripts that do not match English or Bangla in the respective settings. We provide our hypothesis and premise NLI template in the Appendix C.1. We also present additional limitations of the multi-lingual NLI model in the Appendix C.2, to foster research on cross-lingual NLI comprising Bangla.
@@ Results
We further examine model performances across on binary and non-binary questions in the selected Reveal subset in Figure 3. We find that the EngLlama model excels in non-binary questions across both languages, outperforming the BenLlama model in both gen_ans and w_cot_gen_ans settings, even in Bangla by a narrow margin. The stronger performance of BenLlama in the aggregate case is largely motivated by binary questions, in which the model obtains accuracy $80\%$. We also find that while CoT steps have an uneven effect on binary questions, they are consistently beneficial for non-binary ones, across both models and languages. This confirms previous findings on the limited effectiveness of CoT in simpler settings by Liu et al. (2024a), and suggests the benefits of CoT generalize even to less-resourced languages.
@@
To conclude our analysis, we conduct a preliminary investigation into how CoT steps influence model answers. We employ ContextCite Cohen-Wang et al. (2024) to attribute the final answer generated by the model to the provided reasoning steps in the w_cot_gen_ans setting using surrogate linear models, an approach similar to LIME (Ribeiro et al., 2016). Figure 4 presents an overview of our results for the two models across both languages. We observe that in most cases, later steps tend to have a larger influence on the model response. Interestingly, we find that both models accord high importance to Bangla language. We leave the exploration of model interpretability in low-resource languages as an interesting direction for future work.
Figure 4: Importance ratio for EngLlama and BenLlama on w_cot_gen_ans reasoning steps.

### v3

@@ 1 Introduction
Large Language Models (LLMs) have demonstrated remarkable versatility across a wide spectrum of natural language processing tasks (radford2019language). A pivotal breakthrough in enhancing their complex reasoning capabilities has been the introduction of Chain-of-Thought (CoT) prompting (wei2022chain), which encourages models to generate intermediate reasoning steps before arriving at final answers, yielding substantial performance improvements (white2024livebench; wang2022self). Despite these advances, the evaluation of LLM reasoning capabilities remains heavily skewed toward high-resource languages, creating significant gaps in our understanding of how these models perform across linguistically diverse contexts. In this work, we focus specifically on the Bangla language, which boasts 268 million speakers and ranks as the sixth most spoken language globally [^2] https://en.wikipedia.org/wiki/List_of_languages_by_total_number_of_speakers, particularly for its computationally challenging morphological richness choudhury2007evolution; das2010automatic. As the native language of Bangladesh and the second most prominent Indo-Aryan language after Hindi ethnologue_2021, Bangla represents a critical case study for cross-lingual reasoning evaluation. The growing technological transformation in densely populated and economically emerging regions where Bangla is spoken bangladesh-tech-strategy underscores the urgent need for developing faithful AI technologies that can enhance social welfare and economic opportunities. While recent work in Bangla focused on simple extractive or multiple-choice question answering ekram-etal-2022-banglarqa; shafayat-etal-2024-benqa; rony2024banglaquadbengaliopendomainquestion, to our knowledge, no datasets with human-validated reasoning steps are available for this language. This lack of resources hinders our ability to assess and improve the reasoning capabilities of LLMs in the Bangla language.
@@
In this work, we address this gap by introducing Reveal-Bangla, a manually translated Bangla version of a subset of the English Reveal dataset, containing annotated multi-step reasoning chains with gold answers. We exploit our resource and its original English counterpart to evaluate the abilities of two small language models—both proficient in Bangla and English, but one predominantly English-centric, and the other mainly Bangla-centric—in exploiting reasoning step to produce the correct answers given a query, following recent work showing how non-English languages can harm reasoning abilities in LLMs (qi2025modelsreasonlanguagecontrolling).
@@
Moreover, recent cross-lingual studies have revealed that generated reasoning chains often exhibit inconsistencies and produce misleading intermediate steps, raising questions about their explanatory reliability (lanham2023measuringfaithfulnesschainofthoughtreasoning; paul-etal-2024-making). To address these concerns, post-hoc attribution techniques have emerged as valuable tools for analyzing models’ internal processes by assigning importance scores to context elements such as retrieved documents (qi-sarti-etal-2024-model; cohen2024contextcite) and reasoning chains (cohenwang2025learning), thereby revealing their contribution to final predictions. We exploit a similar methodology using the ContextCite method (cohen2024contextcite) to examine how reasoning steps contribute to model answers in English and Bangla, highlighting different patterns of importance across the two languages.
@@ 2 Related Work
Large Language Models (LLMs) operate as probabilistic sequence predictors, estimating the likelihood of the next token given previous context vaswani2017attention; radford2019language. For practical application, explicit training on instructions was found to further improve answer quality (sanh2022multitask; wang2022self). Recently, eliciting reasoning from LLMs, e.g. via step-by-step Chain of Thought reasoning (CoT, wei2022chain), was found to further improve the response accuracy for complex queries.
Some popular reasoning datasets in English include StrategyQA (geva2021did), featuring reasoning- and knowledge-intensive yes/no queries; Fermi (kalyan2021much), comprising estimation questions that require numerical answers and a blend of knowledge and reasoning; MuSiQue trivedi2022musique, which includes multi-hop reasoning questions with free-text entity answers, generated from Wikipedia paragraphs; and Sports Understanding srivastava2022beyond, consisting of yes/no questions that demand reasoning about sports players, leagues, and maneuvers. jacovi2024chain combined the aforementioned datasets and human-annotated each LLM-generated step in terms of attribution relative to provided Wikipedia paragraphs and logical coherence in light of previous reasoning steps. The resulting dataset, dubbed Reveal, was used to prompt capable LLMs in a chain-of-thought setting and analyzed using Natural Language Inference (NLI) classifiers to evaluate model-generated responses. In our work, we manually translate a subset of Reveal into Bangla and adopt their setup to evaluate cross-lingual English-Bangla models.
Recently, jin2024impact explored small and large parameterized models, revealing a linear relationship between accuracy and the number of reasoning steps. We conduct a similar analysis, focusing particularly on small language models (SLMs) with a manageable size (1B parameters). SLMs were recently found capable of high-quality answers in RAG setups (huang-etal-2024-less; small-rag-cs), with relevant input information compensating for their limited reasoning abilities. In this work, we use annotated CoTs produced by larger LLMs to investigate whether SLMs can effectively leverage reasoning information in English and Bangla.
@@ 3.1 Data Collection
We start by selecting a subset of the Reveal dataset. [^3] https://huggingface.co/datasets/google/reveal Provided we want to test the ability of SLMs to obtain the correct answer given valid reasoning chains, we focus specifically on examples having all reasoning steps as either logical or fully attributable to the provided Wikipedia paragraphs. Furthermore, among the three models considered by the Reveal authors to generate answers, we decided to choose the two models with the most answers, i.e., Flan-UL2-20B tay2022ul2 and GPT-3 (text-davinci-003, brown2020language). [^4] We do not include Flan-PaLM-540B longpre2023flan due to our limited evaluation resources.
@@ 4 Evaluation
For our evaluation on Reveal and our Bangla variant, we use Llama-3.2-1B-Instruct (or EngLlama) grattafiori2024llama and BanglaLLama-3.2-1b-bangla-alpaca-orca-instruct-v0.0.1 (or BenLlama) zehady2024bongllama. EngLlama is a popular English-centric multilingual SLM, while BenLlama is a Bangla-centric model fine-tuned from EngLlama using Bangla-Alpaca-Orca, a collection of instruction tuning examples including the popular Alpaca and OpenOrca datasets (alpaca; openorca) automatically translated into Bangla. Importantly, despite their different language focus, both models maintain answering capabilities in both English and Bangla, motivating our cross-lingual analysis.
@@
To verify the accuracy of the model-generated final answer against the actual final answer, we choose mDeBERTa-v3-base-xnli-multilingual-nli-2mil7 laurer_less_2022 model as it is the only NLI model that supports both English and Bangla. We consider entailment labels as correct answers and contradict as otherwise. As this NLI model additionally verdicts neutral, authors manually verify the response to classify it as valid or not. Furthermore, as language detection tools such as langdetect nakatani2010langdetect do not support Bangla, we manually assign contradict to answers generated in scripts that do not match English or Bangla in the respective settings. We provide our hypothesis and premise NLI template in the Appendix C.1. We also present additional limitations of the multi-lingual NLI model in the Appendix C.2, to foster research on cross-lingual NLI comprising Bangla.
@@ Results
We further examine model performances across on binary and non-binary questions in the selected Reveal subset in Figure 3. We find that the EngLlama model excels in non-binary questions across both languages, outperforming the BenLlama model in both gen_ans and w_cot_gen_ans settings, even in Bangla by a narrow margin. The stronger performance of BenLlama in the aggregate case is largely motivated by binary questions, in which the model obtains accuracy $80\%$. We also find that while CoT steps have an uneven effect on binary questions, they are consistently beneficial for non-binary ones, across both models and languages. This confirms previous findings on the limited effectiveness of CoT in simpler settings by liu2024mind, and suggests the benefits of CoT generalize even to less-resourced languages.
@@
To conclude our analysis, we conduct a preliminary investigation into how CoT steps influence model answers. We employ ContextCite cohen2024contextcite to attribute the final answer generated by the model to the provided reasoning steps in the w_cot_gen_ans setting using surrogate linear models, an approach similar to LIME (ribeiro-etal-2016-trust). Figure 4 presents an overview of our results for the two models across both languages. We observe that in most cases, later steps tend to have a larger influence on the model response. This suggests that the models place higher emphasis on answer-specific information located in later steps more than on understanding the context provided in earlier steps. This highlights the inherent limitations of these models in context comprehension, which is essential for answering complex questions. Future research could investigate whether this trend holds with larger model sizes. Additionally, we find that both models accord high importance to the Bangla language. We speculate that Bangla’s morphological richness causes to assign larger values across attention layers. We leave the exploration of model interpretability in low-resource languages as an interesting direction for future work.
Figure 4: Importance ratio for EngLlama and BenLlama on w_cot_gen_ans reasoning steps between 4-4 (lowest) and +4+4 (highest).
@@ C.2.3 Labels entailment where the main final answer is missing
Appendix D Google Translation Errors
D.1 Example on American Football Context
Figure 7: Example of Google Translation’s sub-par performance compared to Manual Translation when given the Source English text with the mistake underlined in pink ( ).
D.2 Example on Basketball Context
Figure 8: Example of Google Translation’s sub-par performance on Basketball context with the mistake underlined in pink ( ).